OpenStudy (anonymous):

tan (1/2 cos ^-1 sqrt5/3)=? sqrt is only for 5?

hero (hero):

calculator?

OpenStudy (anonymous):

No...

OpenStudy (ajprincess):

Let \(\cos^-1\large\frac{\sqrt5}{3}\) be \(\alpha\). Then \(\cos\alpha=\large\frac{\sqrt5}{3}\)  \(\tan\alpha=\large\frac{2}{\sqrt5}\) \(\tan\alpha=\large\frac{2\tan\frac{\alpha}{2}}{1-tan^2\frac{\alpha}{2}}\)  Solve for tan(\(\alpha\)/2) that's vat u need to find.